OpenStudy (anonymous):

A construction crew is lengthening a road. The road started with a length of 54 miles, and the crew is adding 2 miles to the road each day. Let L represent the total length of the road (in miles), and let D represent the number of days the crew has worked. Write an equation relating L to D . Then use this equation to find the total length of the road after the crew has worked 31 days.

OpenStudy (anonymous):

right now I have 54+2(31)=L

OpenStudy (anonymous):

not sure if i'm on the right track though

OpenStudy (sdfgsdfgs):

54+2(31)=L is the right ans to the second part of ur prob. wats the eqn ur prob is asking for in the 1st part?

OpenStudy (anonymous):

I think It want me to use the variables in my equation soo would that just be 54+ 2(31D) =L

OpenStudy (anonymous):

L=54+2D

OpenStudy (anonymous):

and the second answer would be 116 right? =)

OpenStudy (anonymous):

CORRECT.

OpenStudy (anonymous):

thank you =)